OpenStudy (anonymous):

Working together, two pumps can drain a certain pool in 4 hours. If it takes the older pump 9 hours to drain the pool by itself, how long will it take the newer pump to drain the pool on its own?

OpenStudy (anonymous):

Do you know the answer?

OpenStudy (anonymous):

?

OpenStudy (gorv):

let two pump one and two pump one in 9 hour empty 1 tank in 1hour it will empty = 1/9 tank let pump 2 empty in x hour similary pump two will empty 1/x tank together in one hour =1/9 +1/x=(x+9)/9x full tank together they empty in 4 hour (x+9)/9x tank empty in = 1 hour 1tank will= 1/((x+9)/9x)=9x/(9+x)=4hour 9x=4*(9+x) 9x=36+4x 5x=36x=36/h hour

OpenStudy (gorv):

sorry 36/5

OpenStudy (anonymous):

A box of crayons costs $1.75, including tax. Mr. Valentino wants to purchase boxes of crayons for his class and has a $25 budget. Write an inequality to solve for the number of boxes of crayon Mr. Valentino can purchase with his budget.